OpenStudy (anonymous):

Please help the table lists data regarding the average salaries of several professional athletes in the years 1991 and 2001. 1991------262,000 2001-----1,360,000 a. use the data points to find a linear function that fits the data b. use the funtion to predict the average salary in 2005 and 2010 a linear function that fits the data is S(x) = (let x = the number of years since 1990 and let S = the average salary x years for 1990) the predicted average salary for 2005 is--? the predicted average salary for 2010 is--? round to the nearest whole number

OpenStudy (anonymous):

S(1)=262,000 S(11)=1,360,000 slope=( 1360000-262000)/(11-1)=109800 point slope form S(x)-262000=109800(x-1) S(x)=109800x+152200 2005 is x=15, 2010 is x=20 S(15)=1,799,200 S(20)=2,348,200

OpenStudy (anonymous):

rsvitale what is the answer to this part please I'm confuse by your repsonse a linear function that fits the data is S(x) =